# Naval Postgraduate School ML machine learning

### 10 Jan 2024 v1

@@ 1 Introduction
TL leverages knowledge gained from large foundation models to enhance performance on downstream tasks. While deterministic embeddings prevail, variational embeddings provide a promising Bayesian alternative. By using variational inference (VI) to infer posterior distributions over latent features, we obtain variational embeddings which capture uncertainty and enable new analyses of transferred representations. However, the use of variational embeddings in transfer learning (TL) remains relatively unexplored despite their ability to capture uncertainty. Uncertainty estimation is crucial for assessing model credibility and identifying unreliable predictions [1, 2, 3, 4]. Specifically, the variance of variational embeddings provides epistemic uncertainty estimates that indicate when models lack knowledge [2, 3]. This benefits building reliable artificial intelligence systems across audio domains.
@@
While extensive work exists on uncertainty quantification and decomposition for classification problems [10, 11, 12], the focus of that work is on the multi-class case, where each sample contains only a single label. Unfortunately, this restriction hinders the application of existing techniques to real-world datasets where multiple labels may be present in each sample [13, 1, 14]. To address this limitation, we adapt the uncertainty decomposition technique from [11] and [12] for use in multi-label classifications and demonstrate its efficacy on AudioSet.
Our main contributions are: (1) we train two different VI implementations ( Monte Carlo (MC) dropout [15] and Flipout [16]) of the popular ResNet-54 architecture on AudioSet and provide the checkpoints for future research; (2) we adapt the uncertainty decomposition technique from Depeweg et al. [11] and Chai [12] for use in multi-label classification problems; (3) we use the adapted decomposition technique to analyze the uncertainty of our pre-trained model on the AudioSet validation set; and (4) we evaluate TL with a focus on uncertainty quality using three publicly available audio datasets (ESC-50, UrbanSound8K, and DCASE2013). To our knowledge, this is the first time that the uncertainty decomposition technique from [11, 12] has been adapted for use in a multi-label classification problem.
@@ 2.1 Transfer learning in audio
Recently, with popular deep learning frameworks like PyTorch offering pre-trained initializations for modern model architectures, TL has become an integral part of modern model development workflows. Building upon this, research efforts have leveraged the large-scale AudioSet dataset to pre-train deep neural networks for enhanced performance on downstream audio tasks [6, 17]. A common TL approach is to directly extract features from a pre-trained model fixed after the initial training. This method transfers general acoustic knowledge to new tasks without updating the model parameters. However, fine-tuning the pre-trained model by allowing parameter updates during training on the downstream data can further improve results by adapting to the task [6]. In this manuscript, we refer to these techniques as "fixed-feature" and "fine-tuned," respectively. Leveraging large pre-trained models via either technique provides significant performance gains across various audio applications [6, 17].
@@ 3 Methodology
3.1 Datasets
Table 1: Description of datasets used to evaluate VI-PANN embeddings in TL
Dataset Num Clips Num Classes Duration Folds
ESC-50 2,000 50 5 sec 5
UrbanSound8K 8,732 10 $\leq{4}$ sec 10
DCASE2013 100 10 30 sec 5
DCASE2013 [9]: A multi-class dataset with 10 classes representing various acoustic scenes and events. It consists of 100 audio samples, each seconds in duration, and is split into 5-folds for cross-validation. DCASE2013 is commonly used to evaluate acoustic scene/event classification in medium duration recordings.
@@ 3.2 Architecture
In order to implement the Flipout model, we utilize the Bayesian-Torch [22] software package. Using Bayesian-Torch, we convert deterministic layers to Bayesian layers. More specifically, linear layers are converted to LinearFlipout layers and Conv2d layers are converted to Conv2dFlipout. These weight layers are initialized using the MOPED methodology described in [26]. We then modify the cross entropy loss function from [6] to a loss function based on the following form of the negative Evidence Lower Bound (ELBO):
@@ 3.3 Uncertainty quantification and decomposition in multi-label classification
In this work, we train and evaluate BDL models on both multi-class and multi-label classification tasks. In the multi-class case (ESC-50, UrbanSound8K, and DCASE2013), we can directly apply the techniques described in Depeweg et al. [11] and Chai [12]. In the multi-label case (AudioSet), however, we must make a slight adaptation to account for the fact that each class is an independent binary classification problem. Following the methodology in [12], we start by calculating the predictive entropy (i.e., total uncertainty). To calculate the predictive entropy, we first calculate the entropy for each class
@@ 3.5 Bayesian deep learning model pre-training
In order to pre-train our models on AudioSet, we adopt the approach and hyper-parameters from [6]. AudioSet acoustic segments are resampled to 32kHz and converted to log-mel spectrograms using a Hamming window of 1024, a hop size of 320, and 64 mel filter banks. Additionally, we remove frequencies above 14kHz and below 50Hz from the samples. We use a batch size of 32, and an Adam optimizer with a learning rate of 0.001. For the MC dropout variants, we use a dropout rate of 0.2 for convolutional layers, and 0.5 for linear layers.
@@ 3.6 Bayesian transfer learning
    • Initialize a deterministic model with our deterministic PANN using MOPED.
@@ 4.1 Foundation model (AudioSet)
Table 2: AudioSet Pre-Training Results
@@ 4.2 Transfer Learning
Table 3: Transfer Learning Results (Accuracy)
@@
Table 4: Baseline Model Performance (Accuracy)–Training Without TL

### 01 Mar 2024 v2

@@ 1 Introduction
TL leverages knowledge gained from large foundation models to enhance performance on downstream tasks. In the audio domain, the feasibility of transfer learning (TL) has been demonstrated through the successful application of TL techniques in numerous applications ranging from music genre classification to heart sound classification [1, 2, 3, 4, 5, 6]. While deterministic embeddings prevail, variational embeddings provide a promising Bayesian alternative. By using variational inference (VI) to infer posterior distributions over latent features, we obtain variational embeddings which capture uncertainty and enable new analyses of transferred representations. However, the use of variational embeddings in TL remains relatively unexplored despite their ability to capture uncertainty. Uncertainty estimation is crucial for assessing model credibility and identifying unreliable predictions [7, 8, 9, 10]. Specifically, the variance of variational embeddings provides epistemic uncertainty estimates that indicate when models lack knowledge [8, 9]. This benefits building reliable artificial intelligence systems across audio domains.
@@
While extensive work exists on uncertainty quantification and decomposition for classification problems [15, 16, 17], the focus of this work is on the multi-class case, where each sample contains only a single label. Unfortunately, this restriction hinders the application of existing techniques to real-world datasets where multiple labels may be present in each sample [18, 7, 19]. To address this limitation, we describe a new uncertainty decomposition method for use in multi-label classification and demonstrate its efficacy on AudioSet.
This manuscript makes several contributions:
  1. 1.
    We modify the popular ResNet-54 architecture used in [1] to create two distinct VI model variants, namely Monte Carlo (MC) dropout [20] and Flipout [21]. This modification enhances the model by providing access to calibrated epistemic (model) uncertainty information, which was lacking in the original model. We pre-train these VI-PANN models on the AudioSet dataset, and share the resulting checkpoints for future research. [^1] https://github.com/marko-orescanin-nps/VI-PANN
  2. 2.
    We derive a method for decomposing uncertainty information in the multi-label classification scenario.
  3. 3.
    We apply the multi-label uncertainty decomposition technique to analyze the uncertainty of VI-PANNs on the AudioSet validation set, providing insights into the model’s performance in real-world scenarios.
  4. 4.
    We systematically evaluate Bayesian TL techniques with a specific emphasis on the quality of uncertainty. This evaluation is conducted using three publicly available audio datasets (ESC-50, UrbanSound8K, and DCASE2013).
  5. 5.
    In the audio pattern recognition domain, we demonstrate, for the first time, the feasibility of transferring calibrated uncertainty information alongside knowledge from upstream tasks. This transfer markedly enhances a model’s capability to excel in downstream tasks, showcasing an innovative approach in leveraging uncertainty for improved model performance.
@@ 2.1 Transfer learning in audio
Recently, with popular deep learning frameworks like PyTorch offering pre-trained initializations for modern model architectures, TL has become an integral part of modern model development workflows. Building upon this, research efforts have leveraged the large-scale AudioSet dataset to pre-train deep neural networks for enhanced performance on downstream audio tasks [1, 5]. A common TL approach is to directly extract features from a pre-trained model fixed after the initial training. This method transfers general acoustic knowledge to new tasks without updating the model parameters. However, fine-tuning the pre-trained model by allowing parameter updates during training on the downstream data can further improve results by adapting to the task [1]. In this manuscript, we refer to these techniques as “fixed-feature” and “fine-tuned,” respectively. Leveraging large pre-trained models via either technique provides significant performance gains across various audio applications [1, 5].
@@ 3.1 Architecture
In order to implement the Flipout model, we utilize the Bayesian-Torch [26] software package. Using Bayesian-Torch, we convert deterministic layers to Bayesian layers. More specifically, linear layers are converted to LinearFlipout layers and Conv2d layers are converted to Conv2dFlipout. These weight layers are initialized using the MOPED methodology described in [30]. In our case, initialization is done by calling the Bayesian-Torch dnn_to_bnn() function with our pre-trained deterministic model and the default moped_delta parameter of 0.5. We then modify the cross entropy loss function from [1] to a loss function based on the following form of the negative Evidence Lower Bound (ELBO):
@@ 3.2 Uncertainty quantification and decomposition in multi-label classification
In this work, we train and evaluate BDL models on both multi-class and multi-label classification tasks. In the multi-class case (ESC-50, UrbanSound8K, and DCASE2013), we can directly apply the techniques described in Depeweg et al. [16] and Chai [17]. In the multi-label case (AudioSet), however, we must modify the multi-class uncertainty decomposition technique to account for the fact that each class is an independent binary classification problem. Following the methodology in [17], we start by calculating the predictive entropy (i.e., total uncertainty). To calculate the predictive entropy, we first calculate the entropy for each class
@@ 3.4 Bayesian deep learning model pre-training
In order to pre-train our models on AudioSet, we adopt the approach and hyperparameters from [1]. Specifically, to standardize and control for the acoustic pre-processing hyperparameters, enabling a direct and meaningful comparison of model performance between our VI-PANNs and the deterministic models detailed in [1], AudioSet acoustic segments are resampled to 32kHz and converted to log-mel spectrograms using a Hamming window of 1024, a hop size of 320, and 64 mel filter banks. Additionally, following the approach in [1], we remove frequencies above 14kHz and below 50Hz from the samples. For additional details on acoustic pre-processing hyperparameter selection, we refer the interested reader to [1]. We use a batch size of 32, and an Adam optimizer with a learning rate of 0.001. For the MC dropout variants, we use a dropout rate of 0.2 for convolutional layers, and 0.5 for linear layers.
@@ 3.5 Bayesian transfer learning
    • Initialize a deterministic model with our deterministic PANN using MOPED (moped_delta = 0.5), as described in 3.1.
@@
For comparison, Table 1 contains a summary of TL model variant, number of parameters, and the number of multiply-accumulate operations (MACs). The presentation is segmented by dataset, as the input feature shape and the number of classes have distinct impacts on the MACs and number of model parameters, respectively.
Table 1: Model parameter counts and multiply-accumulate operations (MACs) of the three VI model variants used in the transfer learning experiments
Dataset Model Params MACs (per-sample)
ESC-50 Drop 103,340,786 13,536,497,600
Det-Flip 107,639,588 13,545,306,248
Flip 206,620,004 26,653,313,032
UrbanSound8k Drop 103,258,826 13,536,333,760
Det-Flip 107,475,668 13,544,814,568
Flip 206,456,084 26,652,821,352
DCASE2013 Drop 103,258,826 81,862,378,176
Det-Flip 107,475,668 81,870,858,984
Flip 206,456,084 160,695,176,040
@@ 4 Results and Discussion
4.1 Datasets
Table 2: Characteristics of the datasets used to evaluate VI-PANN embeddings in transfer learning
Dataset Num Clips Num Classes Duration Folds
ESC-50 2,000 50 5 sec 5
UrbanSound8K 8,732 10 $\leq{4}$ sec 10
DCASE2013 100 10 30 sec 5
DCASE2013 [14]: A multi-class dataset with 10 classes representing various acoustic scenes and events. It consists of 100 audio samples, each 30 seconds in duration, and is split into 5-folds for cross-validation. DCASE2013 is commonly used to evaluate acoustic scene/event classification in medium duration recordings.
@@ 4.2 Foundation model (AudioSet)
Table 3: Model mean average precision (mAP), area under the receiver operating characteristic curve (AUC), and d-prime after pre-training on the AudioSet dataset
@@ 4.3 Transfer Learning
For reference, we present the results alongside the state-of-the-art (SOTA) results for each dataset. It is essential to clarify that the primary aim of this study was not to achieve SOTA performance on these datasets. Instead, our goal was to demonstrate performance comparable to existing methods while also offering calibrated epistemic uncertainty information. Nonetheless, our VI-PANNs demonstrate comparable performance (within 2 to 3% accuracy) to these SOTA approaches. It’s worth noting that many of the model architectures employed to achieve SOTA performance on these datasets are Transformer-based. In contrast to our approach, these Transformer-based architectures are deterministic and do not provide access to calibrated epistemic uncertainty information.
Table 4: Transfer learning experiment model accuracies on ESC-50, UrbanSound8k, and DCASE2013
@@
SOTA 0.984 [35] 0.908 [36] 0.916 [5]
Table 5: Baseline model accuracies after training on ESC-50, UrbanSound8k, and DCASE2013 without transfer learning